# AutoMMLab: Automatically Generating Deployable Models from Language Instructions for Computer Vision Tasks

### v1

@@ 4.2 Request Understanding
Request understanding stage aims to automatically and accurately understand the user’s request and parse it into a JSON-format configuration.
GPT-assisted Training Data Generation. Tuning LLMs typically require a large amount training data [41, 40], which is exceptionally costly and time-consuming to collect with human crowd-scouring. Inspired by the success of recent LLMs in data generation [65], we leverage ChatGPT/GPT-4 to build our data generation pipeline. For each task type, (i.e. image classification, object detection, semantic segmentation, and keypoint estimation), we first ask 5 professional computer vision researchers to manually design 100 diverse request-config pairs. These human annotations are used as seed examples in in-context-learning to query GPT. In our pipeline, we first design prompts to ask GPT to generate diverse requests, and then use carefully designed prompts to query GPT to parse the requests and generate corresponding configurations. In our preliminary experiments, we ablated the use of ChatGPT and GPT-4 and empirically found that GPT-4 can consistently produce higher quality data. Therefore, we use GPT-4 in our data generation pipeline and generate 1,000 request-config pairs. Finally, the generated data samples are carefully checked and corrected by a group of professional annotators.
RU-LLaMA. Then we use the training data we built to train our request understanding model. Following the mainstream methods of fine-tuning large language models in specific domains, we use LoRA [31] technology to fine-tune the LLaMA2-7B [62] model and obtained our request understanding model RU-LLaMA. We also consider human-in-loop in the request understanding stage. Considering that the “model training requests” requested by users may not be clear and may have missing items, we designed the multi-round request understanding. The keys in the JSON-format configuration we defined can be divided into necessary keys and unnecessary keys. When some keys are missing from the user request, default settings will be adopted for the unnecessary keys (e.g. When the user does not constrain “model”, the best performing model will be selected from the model zoo by default.), and the user will be asked for necessary keys (e.g. If the user does not specify the “object” to be recognized, our system will ask the users).
4.3 Data Selection
Data selection stage aims to use the JSON-format config parsed by RU-LLaMA to automatically retrieve and build request-related training dataset from the dataset zoo.
Figure 4: Examples of (a) data card and (b) model card.
Figure 5: Overview of HPO-LLaMA. At the initial step ($t=1$), HPO-LLaMA proposes a hyperparameter configuration based the model and task description. Model training is then executed using the hyperparameter configuration. The quality of the configuration is measured through model accuracy, and the results are passed back to HPO-LLaMA via a natural language prompt for further iterations ($t>1$).
Dataset Zoo. To facilitate dataset preparation, we construct a comprehensive dataset zoo encompassing a wide variety of public datasets [36, 37, 15, 76]. Each dataset is complemented by a descriptive data card that contains the meta information appertaining to the dataset intricately. Figure 4 (a) presents an illustration of a data card. We take into account AI safety considerations, by ensuring openness, transparency, and reliability of data.
Dataset Selection Pipeline. We design an elaborate pipeline to automatically retrieve pertinent images from the dataset zoo. First, we selectively extract the relevant data cards from the dataset zoo based on the specified task category. Next, if the user requests the specific dataset, we compute the fuzzy match scores, quantifying the similarity between the user-designated name and the name of data cards. These cards are subsequently reordered in a descending pattern, according to their scores. Then the images that meet the requirements are sequentially collected from each dataset. If the annotated categories of images in dataset is the synonyms or hyponyms of the target objects, the images are collected. Synonyms and hyponyms determination is facilitated by the NLTK (Natural Language Toolkit) [3]. Finally, all the collected images constitute the training dataset.
4.4 Model Selection
Model selection automates the process of choosing the best model from the model zoo, based on user requirements.
Model Zoo. We first construct a comprehensive model zoo where each model is accompanied by a detailed model card and the pre-trained model weights. Each model card incorporates attributes such as the model’s name, structure (e.g. network depth), parameters, floating point operations per second (FLOPs), inference speed, and performance. Figure 4 (b) provides a example of a model card. We have considered AI safety issues when constructing the model zoo. We try to keep the model source open and transparent, and all models are pre-trained with reliable public data.
Model Selection Pipeline. We also design a elaborate pipeline able to automatically determine the most suitable model in model zoo, tailored to the user’s model-specific requirements. First, we extract the task-related model cards from the model zoo, predicated on the task categorization specified in parsed config. If the user has preference to specific model, then calculate the fuzzy matching score between the name containing the model structure information in each model card and the user’s specified requirements. Next, we filter out model cards that did not meet the constraints imposed by users regarding the model’s parameters, FLOPs, or the inference speed. Finally, we select the optimal model from the remaining model cards based on fuzzy matching scores and model performance.
4.5 Model Training with Hyperparameter Optimization (HPO)
Hyperparameter optimization (HPO) or hyperparameter tuning is the problem of choosing a set of optimal hyperparameters for model training, which is crucial for the enhancement of the model performance. Hyperparameters are configuration settings that are not learned during training, but are predefined to control the learning process (e.g. learning rate and batch size).
HPO-LLaMA. We propose to tackle the challenging HPO task by training a “black box” optimiser, called HPO-LLaMA. Figure 5 demonstrates an overview of our proposed HPO-LLaMA workflow. Initially, HPO-LLaMA begins by receiving a universally applicable specification of the HPO problem in conjunction with request-specific details relating to the data and model. Following this, it generates a specific hyperparameter setting tailored to the request. Subsequently, HPO-LLaMA acquires the outcomes from the model trainings conducted using the specified hyperparameter setting on the test data and then endeavors to recommend an optimally enhanced hyperparameter setting.
Search Space. We carefully constructed a compact yet expressive search space for HPO, which includes: optimizer type, initial learning rate, learning rate decay policy, weight decay value, batch size, and the training iterations quantity. Considering the heterogeneity of the four tasks, a bespoke search space is delineated for each individual task.
Training HPO-LLaMA. We first adopt GPT-assisted data generation method (introduced in Section 4.2) to generate 100 diverse training requests for each of the four CV tasks. And for each training request, we stochastically select 20 distinct hyperparamter settings from the search space. Consequently, we extensively execute 8,000 model training experiments with various hyperparameter settings. The outcomes of these experiments (model training logs) are compiled into a substantial dataset, comprising triplets of ”request-hyperparameter-performance”. we further conduct a comprehensive evaluation to elucidate the correlation between hyperparameters and model performance. This is primarily to substantiate the appropriateness of the predetermined selection for hyperparameter space. Please refer to the supplementary for more details.
We believe that the data of the triplets contains rich knowledge about model training and hyperparameter tuning. We utilize triplet data to construct both single-round and multi-round dialogues, thereby empowering HPO-LLaMA to activate the capability of predicting efficient hyperparameters from these dialogues. Specifically, when constructing the first round of dialogue, the contextual prompts includes the hyperparameter search space, along with a detailed description of the data and the model. The hyperparameter settings to be predicted are derived from the top-$5$ optimal hyperparameter settings corresponding to the request. In multiple rounds of conversations, the contextual prompt provided to HPO-LLaMA includes historical prompt of previous conversations and the performance of the model trained on the test set for the last time the model predicted hyperparameters. The predicted hyperparameter settings are derived from top-$5$ of all hyperparameter settings corresponding to the request and are the best performing hyperparameter setting in current multi-round dialogue. Finally, we construct the training dataset for up to three dialogue rounds and selectively sample a subset from the entire pool of potential multi-turn conversation data.
We then use LoRA to fine-tune LLaMA-7B and get our HPO-LLaMA. Considering the limited volume of data in a single-round dialogue, we mix dialogue rounds ranging from one to three together to train the HPO-LLaMA.
Table 1: Evaluation of request understanding (RU). Best results are marked in bold. The assessment is based on parsing accuracy.
@@ 4.5 Model Training with Hyperparameter Optimization (HPO)
Mistral-7B - - - -
Llama2-13B - - - -
PaLM 94.11 85.63 92.22 38.75
GPT-3.5-turbo 96.07 93.75 95.56 63.75
GPT-4 98.39 98.13 98.33 86.25
RU-LLaMA 98.57 98.13 98.47 88.75
4.6 Model Deployment
Model deployment in computer vision means integrating a trained computer vision model into a real-world production environment. Model deployment mainly involves model conversion, graph structure optimization, and model quantization. Our AutoMMLab system is equipped with MMDeploy [13], a comprehensive model deployment toolbox, for one-click end-to-end model deployment. It supports a growing list of models that are guaranteed to be exportable to various backends (e.g. ONNX, NCNN, OpenVINO). The model converters enable converting OpenMMLab models into backend models that can be run on target devices. The resulting package is also with a inference SDK, supporting data preprocessing, model forward and postprocessing modules during the model inference phase.
5 Experiments
5.1 Request Understanding
Baseline Models. We built several baseline methods by giving LLMs carefully designed prompts. The prompts comprise an instruction and, optionally, a few demonstrations of the anticipated behavior. We opt to employ both open-source models (Mistral-7B [33] and Llama2-13B [62]) and closed-source models (PaLM [10], GPT-3.5-turbo [4] and GPT-4 [1]) through API calls as baseline models to evaluate the request understanding ability.
Main Results. Table 1 presents the outcomes of various methods evaluated on LAMP utilizing the evaluation metrics delineated in Section 3.1. Without fine-tuning, Mistral-7B and Llama2-13B are unable to parse requests into JSON-format config that conform to predefined formats, scoring 0 in all tasks. RU-LLaMA demonstrates exciting capability in understanding requests and outperforms PaLM, GPT-3.5-turbo and GPT-4 with a 7B-parameter model for both key-level and req-level evaluations.
@@ 5.2 Hyper-parameter Optimization (HPO)
Figure 6: Results of HPO-LLaMA on four tasks: (a) image classification, (b) object detection, (c) semantic segmentation and (d) keypoint detection. Our proposed HPO-LLaMA demonstrates significantly higher efficiency than random sampling baselines.
Random Sampling Baseline. In the realm of HPO, random sampling is usually employed as an important baseline. Specifically, we uniformly sample values from our pre-defined hyperparameter search space. We perform 10 sampling iterations and subsequently identify the optimal configuration to serve as our baselines. Nevertheless, we observed considerable variance across individual trials resulting from this sampling approach. To address this, we repeated the 10-iteration process 1,000 times and reported the average and standard deviation.
LLM baselines. We further compare with several popular LLMs including GPT-3.5-turbo, and GPT-4. Specifically, we used the same prompt as our HPO-LLaMA except that we additionally provided some few-shot examples. As a sanity check, we also build a simple baseline by directly prompting LLaMA2-7B and LLaMA2-13B to solve HPO.
Main Results. Figure 6 shows that our proposed HPO-LLaMA demonstrates a remarkable ability for hyperparameter optimization. It significantly outperforms the random sampling baseline on all four representative tasks. It is worth noting that with only one single iteration trial, our HPO-LLaMA already produces good hyper-parameters.
In Table 2, we compare our proposed HPO-LLaMA with a number of popular LLMs with the 1-iteration evaluation protocol and demonstrate the effectiveness of our proposed HPO-LLaMA. We empirically find that the prompt-engineered LLaMA2 (7B or 13B) lacks strong instruction-following capability and is unable to generate appropriate values in our required data format. An intriguing finding is that after meticulously designing prompts for GPT-3.5-turbo and GPT-4, they exhibit commendable abilities in proposing hyperparameter settings in tasks relating to image classification, object detection, and semantic segmentation. However, in contrast, for the relatively uncommon task of keypoint detection, they fail to suggest useful hyperparameter settings. The HPO-LLaMA model proposed in our study demonstrates the capacity to provide useful hyperparameter recommendations across all four computer vision tasks, and achieves the best performance. At the same time, as the number of iterations increases, HPO-LLaMA can continue to search for better hyperparameter settings.
Table 2: Evaluation of hyper-parameter optimization (HPO). Best results are marked in bold, second best results are underlined.
Model #Iter Cls. Det. Seg. Kpt.
LLaMA2-7B/13B 1 - - - -
GPT-3.5-turbo 1 0.8492 0.3639 0.8522 0.2038
GPT-4 1 0.8610 0.4344 0.8027 0.0958
HPO-LLaMA 1 0.9750 0.4353 0.8538 0.7278
HPO-LLaMA 3 0.9827 0.4397 0.8555 0.7384
@@ 5.3 End-to-end Evaluation
Table 3: End-to-end evaluation on LAMP benchmark. The assessment is based on a grading system scoring from ‘0’ to ‘2’, where ‘0’ denotes ‘total failure’, ‘1’ denotes ‘workable model’, 2 denotes ‘perfect model’. The full score for each task is 40, and the total full score is 160. Best results are marked in bold.
RU model HPO model Cls. Det. Seg. Kpt. Total
PaLM PaLM 16 25 21 18 80
GPT-3.5 GPT-3.5 19 23 26 11 79
GPT-4 GPT-4 24 28 27 14 93
RU-LLaMA HPO-LLaMA 34 33 34 29 130
Table 3 reports the results of end-to-end evaluation on LAMP benchmark. There are four types of model training tasks, i.e. image classification (Cls.), object detection (Det.), semantic segmentation (Seg.) and keypoint detection (Kpt.). We find that our AutoMMLab empowered by RU-LLaMA and HPO-LLaMA significantly outperform other baselines, scoring 130/160 in the overall evaluation and achieves the best performance on all four computer vision tasks.
@@ 6 Conclusion
This paper introduces AutoMMLab, an innovative AutoML system that leverages LLMs to automate the complete process of model production for computer vision tasks using textual instructions. We develop RU-LLaMA for request understanding and propose HPO-LLaMA, a novel approach for effective hyperparameter optimization. Additionally, we present the LAMP (Language-instructed Automated Model Production) Benchmark, which utilizes AutoMMLab to evaluate recent LLMs in the field of AutoML. We hope our work could have broad applicability and inspire further research in related areas.
7 Broad impact and ethics statement
The accessibility of any powerful technology to the general public raises ethical considerations [66], particularly concerning the potential misuse by malicious actors. The deployment of AutoMMLab poses a risk of generating models that may produce toxic or offensive content when exposed to harmful input. To address these ethical concern, our design emphasizes transparency throughout the entire process and incorporates rigorous security measures in the construction of dataset zoo and model zoo, ensuring the reliability of the generated models.
Acknowledgement. We would like to thank Lumin Xu and Kang Chen for valuable discussions.
Appendix A1 Example of End-to-end Model Production
Figure A1: Example of end-to-end model production workflow.
@@ 7 Broad impact and ethics statement
A2.1 Request Generation Prompts
We carefully design prompts for GPT4 and subsequently refine the prompts iteratively to generate requests in batches. Table A1 illustrates the prompt template we designed, which mainly consists of basic prompts, constraints on generated content and anticipated examples. Table A2 shows the constraints we designed.
@@
Table A3: Examples of request.
A2.2 Request Parsing Prompts
Table A4 shows the prompt template we designed for demand analysis, which consists of three parts: basic prompt, JSON format definition of the parsed configuration, and expected examples. Listing LABEL:list:config_json_format shows the JSON format of the configuration we defined. Table A5 presents an example of request parsing. This prompts template is used in multiple places, including: using GPT-4 to generate training data for request understanding, building the baseline for request understanding, and serving as context information when fine-tuning RU-LLaMA.
@@
For the user’s data requirements, you need to summarize: the user’s application scenario, the target object in the image, the data modality, and the dataset required by the user, etc.
For the user’s model requirements, you need to summarize: the tasks that the user wants to achieve, the model that the user specifies to use, the running speed of the model, the number of parameters of the model, the amount of calculation of the model, the measurement method and target value of the model performance.
You need to pay attention that for tasks that require both detection and classification, they need to be divided into detection tasks.
@@
We employ a variety of publicly available datasets to constitute the dataset zoo, designed to accommodate numerous tasks and diverse scenarios. For example, ImageNet1k [36] hosts more than 10 million images spanning across 1,000 categories like animals, plants and vehicles; COCO [37] embraces over 200,000 images and encapsulating location annotations of 80 frequently encountered categories in daily life; Cityscapes [15] holds more than 5,000 images supplied with pixel-level annotations for pedestrians, vehicles, roads, and buildings in urban street scenes; AP10K [76] has 10,000 images with a variety of animal key point annotations. Our dataset zoo supports the continuous expansion of the datasets list, facilitating its applicability to a wider range of tasks and application scenarios.
@@
Table A6 delineates the models currently supported by our model zoo and the corresponding pretrained weights. Our model zoo also supports a growing list of pretrained models for different tasks.
Table A6: Models included in model zoo and corresponding pretrained weights.
Task
Models
Pretrained Weights
Classification
VGG [55], Inception [58], ResNet [24], SEResNet [32], EfficientNet [17], MobileNet [29], MobileViT [46], ShuffleNet [78, 45], ViT [16], SwinTransformer [43]
ImageNet [36]
Objection Detection
FSAF [81], FCOS [60], SSD [42], RetinaNet [38], Faster-RCNN [50], YOLOv3 [49], YOLOX [20], DETR [6]
COCO [37]
Senmantic Segmentation
FCN [53], PSPNet [79], DeepLab [8, 9], UNet [51], APCNet [23], BiSeNet [73, 74], FastFCN [68], UperNet [70], DANet [19], Segformer [71]
Cityscapes [15]
Keypoint Detection
SimpleBaseline2D [69], HRNet [56], LiteHRNet [75]
COCO [37], AP10K [76]
@@
A5.1 Search Space
Table A7 details the hyperparameter search space we set for four computer vision tasks. We further analyze the correlation between the hyperparameters and model performance in the 8000 trained models. For numerical hyperparameters, we use Pearson correlation coefficients to quantitatively measure them. And for categorical hyperparameters, we use boxplots for qualitative analysis. The results are shown in the Table A8, Figure A2, and Figure A3. It can be seen from the charts that the performance of the model has larger correlation with the initial learning rate and optimizer type. Models trained with the AdamW optimizer are more likely to perform better on four computer vision tasks.
@@
Classification
Object detection
Semantic segmentation
Keypoint detection
Optimizer types
SGD [57], Adam [35], AdamW [44], RMSprop [21]
Learning rate decay strategies
MultiStepLR, ConsineAnnealingLR, StepLR, PolyLR
Initial learning rate range
$[10^{-8},0.1]$
Weight decay range
$[10^{-5},0.1]$
Training iteration range
[2000,5000]
[4000,9000]
[2000,7000]
[2000,5000]
Batch size range
[1,64]
[1,16]
[2,8]
[2,64]
Table A8: Correlation between numerical hyperparameters and the final performance of the model. The assessment is based on Pearson correlation coefficient scoring from ‘-1’ to ‘1’, where ‘-1’ denotes ‘complete negative correlation’, ‘1’ denotes ‘complete positive correlation’, and ‘0’ denotes ‘no correlation’.
Classification
Object detection
Semantic segmentation
Keypoint detection
Initial learning rate
-0.22
-0.49
0.17
-0.43
Weight decay
-0.11
0.01
-0.14
-0.07
Training iteration
0.09
-0.01
0.10
-0.01
Batch size
0.08
0.12
0.01
-0.01
@@
38}
A5.2 HPO Prompts
Table A9 presents the template of our hyperparameter optimization prompts. Listing LABEL:list:hpo_json_format shows the hyperparameter search space described in JSON format. Listing LABEL:list:hpo_example_datamodel shows the information of data and model described in JSON format. Listing LABEL:list:hpo_example_hp shows the hyperparameter setting described in JSON format. The HPO prompts serve as the hints in the development of the baseline for the hyperparameter optimization model employing Generative Pre-trained Transformers (GPTs). Additionally, they function as contextual inputs during the training process of HPO-LLaMA.
### Round 1 ###
{{Hyperparameter search space definition in JSON format}}
User:{{Data and model description in JSON format}}
Assistant:{{Hyperparameter setting in JSON format}}
### Round 2 ###
Assistant: {{Hyperparameter setting in JSON format}}
### Round 3 ###
Assistant: {{Hyperparameter setting in JSON format}}
Table A9: Hyperparameter optimization prompts template.
@@
A5.3 Full Experimental Results of Random Sampling
Table A10 presents the comprehensive quantitative results of hyperparameter optimization with the random sampling baseline. The evaluation metrics employed for image classification, object detection, semantic segmentation, and keypoint detection are accuracy, mean Average Precision (mAP), mean Intersection over Union (mIoU), and mAP, respectively.

### v2

@@ 6.1 Request Understanding
Baseline Models. We built several baseline methods by giving LLMs carefully designed prompts. The prompts comprise an instruction and, optionally, a few demonstrations of the anticipated behavior. LLaMA2-7B without instruction fine-tuning cannot accurately follows instructions, so we choose LLaMA2-7B-Chat with instruction fine-tuning as our baseline. As GPTs are currently the most successful LLMs for general language tasks, we opt to employ GPT-3.5-turbo and GPT-4 through API calls as baseline models to evaluate the request understanding ability. We also evaluate the request understanding ability of PaLM2.
Main Results. Table 1 presents the outcomes of various methods evaluated on LAMP utilizing the evaluation metrics delineated in the section of LAMP Benchmark. The findings manifest that LLaMA2-7B-Chat, which did not use the dataset we constructed for fine-tuning, is not able to completely correctly parse the request. Both RU-LLaMA and GPT4 demonstrate exciting capability in understanding requests. RU-LLaMA outperforms PaLM2, GPT-3.5-turbo, and GPT-4 with a 7B-parameter model for both key-level and req-level evaluations.
(a) Image Classification
(b) Object Detection
(c) Semantic Segmentation
(d) Keypoint Detection
Figure 5: HPO results of HPO-LLaMA and random sampling baselines on four tasks : (a) image classification, (b) object detection, (c) semantic segmentation and (d) keypoint detection. HPO-LLaMA demonstrates significantly higher efficiency.
@@
LLaMA2-7B-Chat 85.71 50.00 77.78 0
PaLM2 96.79 88.13 94.86 63.75
GPT-3.5-turbo 96.43 95.63 96.25 72.50
GPT-4 97.50 93.13 96.53 80.00
RU-LLaMA 98.57 96.88 98.20 86.25
Table 1: Evaluation of request understanding (RU). Best results are marked in bold.
Model #R Cls. Det. Seg. Kpt.
BayesianRF 5 0.618$\pm$0.287 0.291$\pm$0.211 0.847$\pm$0.044 0.069$\pm$0.136
BayesianGP 5 0.761$\pm$0.264 0.280$\pm$0.208 0.848$\pm$0.041 0.081$\pm$0.150
LLaMA2-7B 1 0.839$\pm$0.213 0.128$\pm$0.164 0.291$\pm$0.409 0$\pm$0
PaLM2 1 0.964$\pm$0.056 0.367$\pm$0.196 0.845$\pm$0.067 0.719$\pm$0.079
GPT-3.5-turbo 1 0.849$\pm$0.214 0.364$\pm$0.194 0.852$\pm$0.044 0.204$\pm$0.160
GPT-4 1 0.861$\pm$0.188 0.434$\pm$0.147 0.803$\pm$0.194 0.096$\pm$0.158
HPO-LLaMA 1 0.975$\pm$0.028 0.435$\pm$0.148 0.854$\pm$0.042 0.728$\pm$0.051
HPO-LLaMA 3 0.983$\pm$0.020 0.440$\pm$0.150 0.856$\pm$0.043 0.738$\pm$0.053
Table 2: Evaluation of HPO. #R means the number of iteration rounds. The mean and standard deviation on for task are exhibited. Best results are marked in bold, second best results are underlined.
@@ 6.2 Hyper-parameter Optimization (HPO)
Baselines: (1) Random Sampling. In the realm of HPO, random sampling is usually employed as an important baseline. Specifically, we uniformly sample values from our pre-defined hyperparameter search space. We perform sampling for 10 rounds and subsequently identify the optimal configuration to serve as our baselines. Nevertheless, we observed considerable variance across individual trials resulting from this sampling approach. To address this, we repeated the 10-round process 1,000 times and reported the average and standard deviation. (2) Bayesian. We compare with two traditional Bayesian HPO methods, using random forest (BayesianRF) and Gaussian process (BayesianGP) as surrogate models. For each method we perform five rounds of optimization. (3) LLM prompting. We compare with several popular LLMs including LLaMA2-7B-Chat, PaLM2, GPT-3.5-turbo, and GPT-4. Specifically, we directly prompt these general-purpose LLMs with some few-shot examples to produce hyperparameter settings. Note that we used the same prompt as our HPO-LLaMA for fair comparisons.
Main Results. Figure 5 shows that our proposed HPO-LLaMA demonstrates a remarkable ability for hyperparameter optimization. It significantly outperforms the random sampling baseline on all four representative tasks. It is worth noting that with only one single round trial, our HPO-LLaMA already produces good hyper-parameters. In Table 2, we compare our proposed HPO-LLaMA with Bayesian HPO methods using a 5-round evaluation protocol, as well as several popular LLMs with a 1-round evaluation protocol, demonstrating the effectiveness of our approach. In our experiment, we observed that relying on the few-shot capability of LLaMA2-7B-Chat to suggest hyperparameters doesn’t consistently yield the anticipated output, and the effectiveness of the suggested hyperparameters is poor. An intriguing finding is that after meticulously designing prompts for GPT-3.5-turbo and GPT-4, they exhibit commendable abilities in proposing hyperparameter settings in tasks relating to image classification, object detection, and semantic segmentation. However, in contrast, for the relatively uncommon task of keypoint detection, they fail to suggest useful hyperparameter settings. The HPO-LLaMA model proposed in our study demonstrates the capacity to provide useful hyperparameter recommendations across all four computer vision tasks, and achieves the best performance. At the same time, as the number of rounds increases, HPO-LLaMA can continue to search for better hyperparameter settings.
RU HPO Cls. Det. Seg. Kpt. Total
LLaMA2-7B LLaMA2-7B 0 0 0 0 0
PaLM2 PaLM2 14 25 27 15 81
GPT-3.5-turbo GPT-3.5-turbo 24 24 25 11 84
GPT-4 GPT-4 17 27 29 14 87
RU-LLaMA HPO-LLaMA 31 31 32 18 112
Table 3: End-to-end evaluation on LAMP benchmark. The assessment is based on a grading system scoring from ‘0’ to ‘2’, where ‘0’ denotes ‘total failure’, ‘1’ denotes ‘workable model’, 2 denotes ‘perfect model’. The full score for each task is 40, and the total full score is 160.
@@ 6.3 End-to-end Evaluation
Table 3 reports the results of end-to-end evaluation on LAMP benchmark. There are four types of model training tasks, i.e. image classification (Cls.), object detection (Det.), semantic segmentation (Seg.) and keypoint detection (Kpt.). We employ various LLMs for request understanding (RU) and hyperparameter optimization (HPO) respectively. We find that our AutoMMLab empowered by RU-LLaMA and HPO-LLaMA significantly outperform GPT-3.5 and GPT-4 with a 7B-parameter model, scoring 112/160 in the overall evaluation and achieves the best performance on all four computer vision tasks.
@@ 7 Conclusion
This paper presents AutoMMLab, an innovative AutoML system that harnesses the power of LLMs to fully automate the model development process for computer vision tasks, guided by natural language instructions. We develop RU-LLaMA for understanding the user’s requests and scheduling the whole workflow. Furthermore, we propose HPO-LLaMA, a novel LLM-empowered method for effective hyperparameter optimization. Additionally, we present the LAMP (Language-instructed Automated Model Production) Benchmark to facilitate the evaluation of end-to-end prompt-based model production. We equip our AutoMMLab with various LLMs and test on LAMP, demonstrating the superiority of our proposed RU-LLaMA and HPO-LLaMA. We hope our work could have broad applicability and inspire further research in related areas.
@@
A1.1 Request Generation Prompts
We carefully design prompts for GPT4 and subsequently refine the prompts iteratively to generate requests in batches. Table A1 illustrates the prompt template we designed, which mainly consists of basic prompts, constraints on generated content and anticipated examples. Table A3 shows the constraints we designed. Table A2 exhibits three examples of generated requests.
@@
Table A4: Example of request parsing.
A1.2 Request Parsing Prompts
Table A5 shows the prompt template we designed for request parsing, which consists of three parts: basic prompt, JSON format definition of the parsed configuration, and expected examples. Listing 1 shows the JSON format of the configuration we defined. Table A4 presents an example of request parsing. This prompts template is used in multiple places, including: using GPT-4 to generate training data for request understanding, building the baseline for request understanding, and serving as context information when fine-tuning RU-LLaMA.
@@
For the user’s data requirements, you need to summarize: the user’s application scenario, the target object in the image, the data modality, and the dataset required by the user, etc. For the user’s model requirements, you need to summarize: the tasks that the user wants to achieve, the model that the user specifies to use, the running speed of the model, the number of parameters of the model, the amount of calculation of the model, the measurement method and target value of the model performance. For the user’s deployment requirements, you need to summarize: information such as the deployment environment and equipment required by the user. You need to pay attention that for tasks that require both detection and classification, they need to be divided into detection tasks.
@@
We employ a variety of publicly available datasets to constitute the dataset zoo, designed to accommodate numerous tasks and diverse scenarios. For example, ImageNet1k (Krizhevsky, Sutskever, and Hinton 2017) hosts more than 10 million images spanning across 1,000 categories like animals, plants and vehicles; COCO (Lin et al. 2014) embraces over 200,000 images and encapsulating location annotations of 80 frequently encountered categories in daily life; Cityscapes (Cordts et al. 2016) holds more than 5,000 images supplied with pixel-level annotations for pedestrians, vehicles, roads, and buildings in urban street scenes; AP10K (Yu et al. 2021c) has 10,000 images with a variety of animal key point annotations. Our dataset zoo supports the continuous expansion of the datasets list, facilitating its applicability to a wider range of tasks and application scenarios.
@@
Table A6 delineates the models currently included by our model zoo. Our model zoo also supports a growing list of pretrained models for different tasks. Regarding image classification, our model zoo contains popular CNN models (e.g. ResNet (He et al. 2016) and EfficientNet (Tan and Le 2019)), Vision Transformers (e.g. ViT (Dosovitskiy et al. 2020)), etc. In terms of object detection, the object detection frameworks contained in our model zoo include: two-stage RCNN series (Girshick 2015; Ren et al. 2015), one-stage YOLO series (Redmon and Farhadi 2018; Ge et al. 2021), DETR-based detectors (Zhu et al. 2021), etc. Concerning semantic segmentation, we have DeepLab (Chen et al. 2017, 2018), UPerNet (Xiao et al. 2018), Segformer (Xie et al. 2021), etc. For pose estimation, we have SimpleBaseline2D (Xiao, Wu, and Wei 2018), HRNet (Sun et al. 2019), LiteHRNet (Yu et al. 2021b), etc.
Table A6: Models included in model zoo.
Task Models
Classification VGG (Simonyan and Zisserman 2014), Inception (Szegedy et al. 2016), ResNet (He et al. 2016), SEResNet (Hu, Shen, and Sun 2018), EfficientNet (Feurer et al. 2015), MobileNet (Howard et al. 2019), MobileViT (Mehta and Rastegari 2021), ShuffleNet (Zhang et al. 2018; Ma et al. 2018), ViT (Dosovitskiy et al. 2021), SwinTransformer (Liu et al. 2021)
Objection Detection FSAF (Zhu, He, and Savvides 2019), FCOS (Tian et al. 2019), SSD (Liu et al. 2016), RetinaNet (Lin et al. 2017), Faster-RCNN (Ren et al. 2017), YOLOv3 (Redmon and Farhadi 2018), YOLOX (Ge et al. 2021), DETR (Carion et al. 2020)
Senmantic Segmentation FCN (Shelhamer, Long, and Darrell 2017), PSPNet (Zhao et al. 2017), DeepLab (Chen et al. 2017, 2018), UNet (Ronneberger, Fischer, and Brox 2015), APCNet (He et al. 2019), BiSeNet (Yu et al. 2018, 2021a), FastFCN (Wu et al. 2019), UperNet (Xiao et al. 2018), DANet (Fu et al. 2019), Segformer (Xie et al. 2021)
Keypoint Detection SimpleBaseline2D (Xiao, Wu, and Wei 2018), HRNet (Sun et al. 2019), LiteHRNet (Yu et al. 2021b)
@@
A4.1 Search Space
Table A8 details the hyperparameter search space we set for four computer vision tasks. We further analyze the correlation between the hyperparameters and model performance in the 8000 trained models. For numerical hyperparameters, we use Pearson correlation coefficients for quantitative analysis. And for categorical hyperparameters, we use boxplots for qualitative analysis. The results are shown in the Table A8, Figure A2, and Figure A2. It can be seen from the charts that the performance of the model has larger correlation with the initial learning rate and optimizer type. Models trained with the AdamW optimizer are more likely to perform better on four computer vision tasks.
@@
Classification Object detection Semantic segmentation Keypoint detection
Optimizer types SGD (Sutskever et al. 2013), Adam (Kingma and Ba 2017), AdamW (Loshchilov and Hutter 2019), RMSprop (Graves 2014)
Learning rate decay strategies MultiStepLR, ConsineAnnealingLR, StepLR, PolyLR
Initial learning rate range $[10^{-8},0.1]$
Weight decay range $[10^{-5},0.1]$
Training iteration range [2000,5000] [4000,9000] [2000,7000] [2000,5000]
Batch size range [1,64] [1,16] [2,8] [2,64]
Table A8: Correlation between numerical hyperparameters and the final performance of the model. The assessment is based on Pearson correlation coefficient scoring from ‘-1’ to ‘1’, where ‘-1’ denotes ‘strong negative correlation’, ‘1’ denotes ‘strong positive correlation’, and ‘0’ denotes ‘no correlation’.
Classification Object detection Semantic segmentation Keypoint detection
Initial learning rate -0.22 -0.49 0.17 -0.43
Weight decay -0.11 0.01 -0.14 -0.07
Training iteration 0.09 -0.01 0.10 -0.01
Batch size 0.08 0.12 0.01 -0.01
@@
A4.2 HPO Prompts
Table A9 presents the template of our hyperparameter optimization prompts. Listing 2 shows the hyperparameter search space described in JSON format. Listing 3 shows the information of data and model described in JSON format. Listing 4 shows the hyperparameter setting described in JSON format. The HPO prompts serve as the hints in the development of the baseline for the hyperparameter optimization model employing Generative Pre-trained Transformers (GPTs). Additionally, they function as contextual inputs during the training process of HPO-LLaMA.
Table A9: Hyperparameter optimization prompts template.
### Round 1 ###
{{Hyperparameter search space definition in JSON format}}
User:{{Data and model description in JSON format}}
Assistant:{{Hyperparameter setting in JSON format}}
### Round 2 ###
Assistant: {{Hyperparameter setting in JSON format}}
### Round 3 ###
Assistant: {{Hyperparameter setting in JSON format}}
@@
38 }
@@
Appendix A5 Details of Experimental Setting and Results